# Wasserstein Convergence of Score-based Generative Models under Semiconvexity and Discontinuous Gradients

### v1

@@ Appendix C Proof of the Main Results
$\begin{split}&\leq(2+3\zeta+4\mathsf{K}_{3}(1+2T^{\alpha}))\int_{0}^{T-\epsilon}\mathbb{E}[|Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}}|^{2}]\ \text{d}t\\ &\quad+\zeta^{-1}\gamma(T-\epsilon)C_{\mathsf{EMose},2}+8\zeta^{-1}\gamma^{2\alpha}(T-\epsilon)\mathsf{K}_{1}^{2}(1+4\mathbb{E}[|\hat{\theta}|^{2}])\\ &\quad+4\zeta^{-1}\mathsf{K}^{2}_{3}(1+2T^{\alpha})^{2}\int_{0}^{T-\epsilon}\mathbb{E}[|\widehat{Y}^{\text{EM}}_{t}-\widehat{Y}^{\text{EM}}_{\lfloor t/\gamma\rfloor\gamma}|^{2}]\mathrm{d}t\\ &\leq(2+3\zeta+4\mathsf{K}_{3}(1+2T^{\alpha}))\int_{0}^{T-\epsilon}\mathbb{E}[|Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}}|^{2}]\ \text{d}t\\ &\quad+\zeta^{-1}\gamma(T-\epsilon)C_{\mathsf{EMose},2}(1+4\mathsf{K}^{2}_{3}(1+2T^{\alpha})^{2})\\ &\quad+8\zeta^{-1}\gamma^{2\alpha}(T-\epsilon)\mathsf{K}_{1}^{2}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})\\ &\leq e^{(2+3\zeta+4\mathsf{K}_{3}(1+2T^{\alpha}))(T-\epsilon)}\\ &\quad\times\Bigg{(}\zeta^{-1}\gamma(T-\epsilon)C_{\mathsf{EMose},2}(1+4\mathsf{K}^{2}_{3}(1+2T^{\alpha})^{2})\\ &\qquad\quad+8\zeta^{-1}\gamma^{2\alpha}(T-\epsilon)\mathsf{K}_{1}^{2}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})\Bigg{)}.\end{split}$ (60)
Using (60) and $t_{K}=T-\epsilon$, we have
@@
$\begin{split}W_{2}(\mathscr{L}(Y_{T-\epsilon}^{\text{aux}}),\mathscr{L}(\widehat{Y}_{T-\epsilon}^{\text{EM}}))&\leq\gamma^{1/2}\zeta^{-1/2}(T-\epsilon)^{1/2}e^{(1+(3/2)\zeta+2\mathsf{K}_{3}(1+2T^{\alpha}))(T-\epsilon)}\\ &\quad\times(C^{1/2}_{\mathsf{EMose},2}(1+2\mathsf{K}_{3}(1+2T^{\alpha}))+2\sqrt{2}\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}).\end{split}$ (61)
Final upper bound on $W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}}).$
Substituting (51), (55), (58), and (61) into (49), we have
@@
$\begin{split}W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})&\leq(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{d})2\sqrt{\epsilon}\\ &\quad+\sqrt{2}(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{d})e^{-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t-\epsilon}\\ &\quad+\sqrt{2\zeta^{-1}}e^{(1+\zeta)(T-\epsilon)-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t}\sqrt{\varepsilon_{\text{SN}}}\\ &\quad+\gamma^{1/2}\zeta^{-1/2}(T-\epsilon)^{1/2}e^{(1+(3/2)\zeta+2\mathsf{K}_{3}(1+2T^{\alpha}))(T-\epsilon)}\\ &\qquad\times(C^{1/2}_{\mathsf{EMose},2}(1+2\mathsf{K}_{3}(1+2T^{\alpha}))+2\sqrt{2}\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}).\end{split}$ (62)
@@
The bound for $W_{2}(\mathscr{L}(\widehat{Y}_{K}^{\text{EM}}),\pi_{\mathsf{D}})$ in (62) can be made arbitrarily small by appropriately choosing parameters including $\epsilon,T,\varepsilon_{\text{SN}}$ and $\gamma$. More precisely, for any $\delta>0$, we first choose $0<\epsilon<\epsilon_{\delta}$ with $\epsilon_{\delta}$ given in Table 2 such that the first term on the right-hand side of (62) is
@@
Next, we choose $T>T_{\delta}$ with $T_{\delta}$ given in Table 2 such that the second term on the right-hand side of (62) is
@@
Next, we turn to the third term on the right-hand side of (62). We choose $0<\varepsilon_{\text{SN}}<\varepsilon_{\text{SN},\delta}$ with $\varepsilon_{\text{SN},\delta}$ given in Table 2 such that
@@
Finally, we choose $0<\gamma<\gamma_{\delta}$ with $\gamma_{\delta}$ given in Table 2 such that the fourth term on the right-hand side of (62) is
@@
Using (63), (64), (65), and (66), we obtain $W_{2}(\mathscr{L}(\widehat{Y}_{K}^{\text{EM}}),\pi_{\mathsf{D}})<\delta$. ∎
Proof of Theorem 3.14.
Using the splitting (49), the proof follows along the same lines of the Proof of Theorem 3.12 for the estimation of the error bounds of the terms $W_{2}(\pi_{\mathsf{D}},\mathscr{L}(Y_{t_{K}}))$, $W_{2}(\mathscr{L}(Y_{t_{K}}),\mathscr{L}(\widetilde{Y}_{t_{K}}))$, and $W_{2}(\mathscr{L}(\widetilde{Y}_{t_{K}}),\mathscr{L}(Y_{t_{K}}^{\text{aux}}))$. The error bound for $W_{2}(\mathscr{L}(Y_{t_{K}}^{\text{aux}}),\mathscr{L}(Y_{K}^{\text{EM}}))$ is derived in [7, Proof of Theorem 10]. Putting these four estimates together leads to (25) and (26). ∎
@@ Appendix D Modified Half-Normal Distribution
In this section, we recall the probability density function of the modified half-normal distribution, see e.g. [59], used in Section 3.4.1 and defined as
@@
is the Fox–Wright function [19, 66]. We point out that the half-normal distribution, truncated normal distribution, gamma distribution, and square root of the gamma distribution are all special cases of the modified Half-Normal distribution (67). The distribution (27) follows by taking the symmetric extension of (67), i.e. $g(|x|)/2$, and choosing $\upsilon=1$ and $\psi=-1$.
Appendix E Tables of Constants
Table 2 displays full expressions for constants which appear in Theorem 3.12 and Theorem 3.14.
Table 2: Explicit expressions for the constants in Theorem 3.12 and Theorem 3.14.
@@ Appendix E Tables of Constants
$C_{\mathsf{EM},2}(T)$ $O(Me^{T^{2\alpha+1}}T^{2\alpha+1}\widetilde{\varepsilon}_{\text{AL}})$ $&e^{T(4+8\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))}\\ &\quad\times(\mathbb{E}[|\widehat{Y}_{0}^{\text{EM}}|^{2}]+16\mathsf{K}^{2}_{\text{Total}}T(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})(1+T^{2\alpha})+2dT)\end{aligned}$
@@
$C_{\mathsf{EM},4}(T)$ $O(d^{2}e^{T^{4\alpha+1}}T^{4\alpha+1})$ $&e^{T(\frac{21}{2}+128\mathsf{K}^{4}_{\text{Total}}(1+T^{4\alpha}))}\\ &\quad\times(\mathbb{E}[|\widehat{Y}_{0}^{\text{EM}}|^{4}]+1024\mathsf{K}^{4}_{\text{Total}}T(1+\mathbb{E}[|\hat{\theta}|^{4}])(1+T^{4\alpha})+8(d^{2}+4d+4)T)\end{aligned}$
@@
$C_{\mathsf{EMose},2}$ $O(de^{T^{2\alpha+1}}T^{4\alpha+1}\widetilde{\varepsilon}_{\text{AL}})$ $2(C_{\mathsf{EM},2}(T)+\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(16C_{\mathsf{EM},2}(T)+32(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})))+2d\end{aligned}$
@@
$C_{4}(T,\epsilon)$ $O(\sqrt{d}e^{T^{2\alpha+1}}T^{3\alpha+1}\widetilde{\varepsilon}^{1/2}_{\text{AL}})$ $&\zeta^{-1/2}(T-\epsilon)^{1/2}e^{(1+(3/2)\zeta+2\mathsf{K}_{3}(1+2T^{\alpha}))(T-\epsilon)}\\ &\qquad\times(C^{1/2}_{\mathsf{EMose},2}(1+2\mathsf{K}_{3}(1+2T^{\alpha}))+2\sqrt{2}\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2})\end{aligned}$
@@
$C_{\mathsf{EMose},4}$ $O(d^{2}e^{T^{4\alpha+1}}T^{8\alpha+1})$ $8(C_{\mathsf{EM},4}(T)+\mathsf{K}^{4}_{\text{Total}}(1+T^{4\alpha})(1024C_{\mathsf{EM},4}(T)+8192(1+\mathbb{E}[|\hat{\theta}|^{4}])))+144d^{2}\end{aligned}$
@@
$\widetilde{C}_{4}(T,\epsilon)$ $O(de^{T^{4\alpha+1}}T^{4\alpha+1}\widetilde{\varepsilon}^{1/4}_{\text{AL}})$ $&\sqrt{2}e^{2(1+\zeta+\mathsf{K}_{3}(1+2T^{\alpha}+4\mathsf{K}_{3}(1+4T^{2\alpha})))(T-\epsilon)}\sqrt{T-\epsilon}\\ &\quad\ \times\Bigg{(}\mathsf{K}_{4}^{2}\zeta^{-1}(1+4T^{2\alpha})C_{\mathsf{EMose},4}+4d(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\quad+2\zeta^{-1}\mathsf{K}_{1}^{2}(1+8(\widetilde{\varepsilon}_{\text{AL}}+|\theta^{*}|^{2}))\\ &\qquad\quad+4\zeta^{-1}d(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\quad\quad\times[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))C_{\mathsf{EM},2}(T)\\ &\qquad\qquad\qquad+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})]\\ &\qquad\quad+2[(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}C^{1/2}_{\mathsf{EMose},2}+2\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}]\\ &\qquad\qquad\times[d\sqrt{2}(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}]\Bigg{)}^{1/2}\end{aligned}$
$\epsilon_{\delta}$ $\delta^{2}/(64(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{d})^{2})$
@@
$T_{\delta}$ Obtained solving $T>T_{\delta}$ using Proposition 3.10, i.e., $&\ln(\mu(e^{2T}-1)+1)+(K/\mu+1)/(\mu(e^{2T}-1)+1)\\ &>\ln(4\sqrt{2}((\mathbb{E}[|X_{0}|^{2}])^{1/2}+\sqrt{d})/\delta)+2\int_{0}^{\epsilon}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t+K/\mu+1-\epsilon\end{aligned}$
@@
$\varepsilon_{\text{SN},\delta}$ $(\delta^{2}\zeta/32)e^{-2(1+\zeta)(T-\epsilon)+4\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t}\end{aligned}$
@@
$\gamma_{\delta}$ $&(\delta^{2}\zeta/16)(T-\epsilon)^{-1}e^{-2(1+(3/2)\zeta+2\mathsf{K}_{3}(1+2T^{\alpha}))(T-\epsilon)}\\ &\quad\times(C^{1/2}_{\mathsf{EMose},2}(1+2\mathsf{K}_{3}(1+2T^{\alpha}))+2\sqrt{2}\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2})^{-2}\end{aligned}$
@@
$\widetilde{\gamma}_{\delta}$ $\min\Bigg{\{}&(\delta/(4\sqrt{2}))^{1/\alpha}(T-\epsilon)^{-1/(2\alpha)}e^{-(2/\alpha)(1+\zeta+\mathsf{K}_{3}(1+2T^{\alpha}+4\mathsf{K}_{3}(1+4T^{2\alpha})))(T-\epsilon)}\\ &\times\Bigg{(}\mathsf{K}_{4}^{2}\zeta^{-1}(1+4T^{2\alpha})C_{\mathsf{EMose},4}+4d(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad+2\zeta^{-1}\mathsf{K}_{1}^{2}(1+8(\widetilde{\varepsilon}_{\text{AL}}+|\theta^{*}|^{2}))\\ &\qquad+4\zeta^{-1}d(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\quad\times[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))C_{\mathsf{EM},2}(T)\\ &\qquad\qquad+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})]\\ &\qquad+2[(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}C^{1/2}_{\mathsf{EMose},2}+2\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}]\\ &\qquad\qquad\times[d\sqrt{2}(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}]\Bigg{)}^{-1/(2\alpha)},1\Bigg{\}}\end{aligned}$

### v2

@@ Appendix C Proof of the Main Results
$\begin{split}\mathbb{E}\left[|Y_{T-\epsilon}^{\text{aux}}-\widehat{Y}_{T-\epsilon}^{\text{EM}}|^{2}\right]&\leq(2+3\zeta)\int_{0}^{T-\epsilon}\mathbb{E}[|Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}}|^{2}]\ \text{d}t+\zeta^{-1}\int_{0}^{T-\epsilon}\mathbb{E}[|\widehat{Y}_{t}^{\text{EM}}-\widehat{Y}^{\text{EM}}_{\lfloor t/\gamma\rfloor\gamma}|^{2}]\ \text{d}t\\ &\quad+4\mathsf{K}_{3}(1+2T^{\alpha})\int_{0}^{T-\epsilon}\mathbb{E}[|Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}}|^{2}]\mathrm{d}t\\ &\quad+2\zeta^{-1}\int_{0}^{T-\epsilon}\mathbb{E}[|s(T-t,\hat{\theta},\widehat{Y}^{\text{EM}}_{t})-s(T-\lfloor t/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor t/\gamma\rfloor\gamma}|^{2}]\mathrm{d}t\\ &\leq(2+3\zeta+4\mathsf{K}_{3}(1+2T^{\alpha}))\int_{0}^{T-\epsilon}\mathbb{E}[|Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}}|^{2}]\ \text{d}t\\ &\quad+\zeta^{-1}\gamma(T-\epsilon)C_{\mathsf{EMose},2}+8\zeta^{-1}\gamma^{2\alpha}(T-\epsilon)\mathsf{K}_{1}^{2}(1+4\mathbb{E}[|\hat{\theta}|^{2}])\\ &\quad+4\zeta^{-1}\mathsf{K}^{2}_{3}(1+2T^{\alpha})^{2}\int_{0}^{T-\epsilon}\mathbb{E}[|\widehat{Y}^{\text{EM}}_{t}-\widehat{Y}^{\text{EM}}_{\lfloor t/\gamma\rfloor\gamma}|^{2}]\mathrm{d}t\\ &\leq(2+3\zeta+4\mathsf{K}_{3}(1+2T^{\alpha}))\int_{0}^{T-\epsilon}\mathbb{E}[|Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}}|^{2}]\ \text{d}t\\ &\quad+\zeta^{-1}\gamma(T-\epsilon)C_{\mathsf{EMose},2}(1+4\mathsf{K}^{2}_{3}(1+2T^{\alpha})^{2})\\ &\quad+8\zeta^{-1}\gamma^{2\alpha}(T-\epsilon)\mathsf{K}_{1}^{2}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})\\ &\leq e^{(2+3\zeta+4\mathsf{K}_{3}(1+2T^{\alpha}))(T-\epsilon)}\\ &\quad\times\Bigg(\zeta^{-1}\gamma(T-\epsilon)C_{\mathsf{EMose},2}(1+4\mathsf{K}^{2}_{3}(1+2T^{\alpha})^{2})\\ &\qquad\quad+8\zeta^{-1}\gamma^{2\alpha}(T-\epsilon)\mathsf{K}_{1}^{2}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})\Bigg).\end{split}$ (62)
Using 62 and $t_{J}=T-\epsilon$, we have
@@
$\begin{split}W_{2}(\mathcal{L}(Y_{T-\epsilon}^{\text{aux}}),\mathcal{L}(\widehat{Y}_{T-\epsilon}^{\text{EM}}))&\leq\gamma^{1/2}\zeta^{-1/2}(T-\epsilon)^{1/2}e^{(1+(3/2)\zeta+2\mathsf{K}_{3}(1+2T^{\alpha}))(T-\epsilon)}\\ &\quad\times(C^{1/2}_{\mathsf{EMose},2}(1+2\mathsf{K}_{3}(1+2T^{\alpha}))+2\sqrt{2}\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}).\end{split}$ (63)
Final upper bound on $W_{2}(\mathcal{L}(Y_{J}^{\text{EM}}),\pi_{\mathsf{D}}).$
Substituting 53, 57, 60, and 63 into 51, we have
@@
$\begin{split}W_{2}(\mathcal{L}(Y_{J}^{\text{EM}}),\pi_{\mathsf{D}})&\leq(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{d})2\sqrt{\epsilon}\\ &\quad+\sqrt{2}(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{d})e^{-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t-\epsilon}\\ &\quad+\sqrt{2\zeta^{-1}}e^{(1+\zeta)(T-\epsilon)-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t}\sqrt{\varepsilon_{\text{SN}}}\\ &\quad+\gamma^{1/2}\zeta^{-1/2}(T-\epsilon)^{1/2}e^{(1+(3/2)\zeta+2\mathsf{K}_{3}(1+2T^{\alpha}))(T-\epsilon)}\\ &\qquad\times(C^{1/2}_{\mathsf{EMose},2}(1+2\mathsf{K}_{3}(1+2T^{\alpha}))+2\sqrt{2}\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}).\end{split}$ (64)
@@
The bound for $W_{2}(\mathcal{L}(\widehat{Y}_{J}^{\text{EM}}),\pi_{\mathsf{D}})$ in 64 can be made arbitrarily small by appropriately choosing parameters including $\epsilon,T,\varepsilon_{\text{SN}}$ and $\gamma$. More precisely, for any $\delta>0$, we first choose $0<\epsilon<\epsilon_{\delta}$ with $\epsilon_{\delta}$ given in Table 3 such that the first term on the right-hand side of 64 is
@@
Next, we choose $T>T_{\delta}$ with $T_{\delta}$ given in Table 3 such that the second term on the right-hand side of 64 is
@@
Next, we turn to the third term on the right-hand side of 64. We choose $0<\varepsilon_{\text{SN}}<\varepsilon_{\text{SN},\delta}$ with $\varepsilon_{\text{SN},\delta}$ given in Table 3 such that
@@
Finally, we choose $0<\gamma<\gamma_{\delta}$ with $\gamma_{\delta}$ given in Table 3 such that the fourth term on the right-hand side of 64 is
@@
Using 65, 66, 67, and 68, we obtain $W_{2}(\mathcal{L}(\widehat{Y}_{J}^{\text{EM}}),\pi_{\mathsf{D}})<\delta$. ∎
Proof of Theorem 21.
Using the splitting 51, the proof follows along the same lines of the Proof of Theorem 19 for the estimation of the error bounds of the terms $W_{2}(\pi_{\mathsf{D}},\mathcal{L}(Y_{t_{J}}))$, $W_{2}(\mathcal{L}(Y_{t_{J}}),\mathcal{L}(\widetilde{Y}_{t_{J}}))$, and $W_{2}(\mathcal{L}(\widetilde{Y}_{t_{J}}),\mathcal{L}(Y_{t_{J}}^{\text{aux}}))$. The error bound for $W_{2}(\mathcal{L}(Y_{t_{J}}^{\text{aux}}),\mathcal{L}(Y_{J}^{\text{EM}}))$ is derived along the same lines of Bruno et al. (2025, Proof of Theorem 10). Putting these four estimates together leads to 27 and 28. ∎
@@ Appendix D Modified Half-Normal Distribution
In this section, we recall the probability density function of the modified half-normal distribution, see e.g., Sun et al. (2023), used in Section 3.4.1 and defined as
@@
is the Fox–Wright function (Fox, 1928; Wright, 1935). We point out that the half-normal distribution, truncated normal distribution, gamma distribution, and square root of the gamma distribution are all special cases of the modified Half-Normal distribution 69. The distribution 29 follows by taking the symmetric extension of 69, i.e. $g(|x|)/2$, and choosing $\upsilon=1$ and $\psi=-1$.
Appendix E Table of Constants
Table 3 displays full expressions for constants which appear in Theorem 19 and Theorem 21.
Table 3: Explicit expressions for the constants in Theorem 19 and Theorem 21.
@@ Appendix E Table of Constants
$C_{\mathsf{EM},2}(T)$ $O(Me^{T^{2\alpha+1}}T^{2\alpha+1}\widetilde{\varepsilon}_{\text{AL}})$ $\begin{aligned} &e^{T(4+8\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))}\\ &\qquad\times(\mathbb{E}[|\widehat{Y}_{0}^{\text{EM}}|^{2}]+16\mathsf{K}^{2}_{\text{Total}}T(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})(1+T^{2\alpha})+2dT)\end{aligned}$
@@
$C_{\mathsf{EM},4}(T)$ $O(d^{2}e^{T^{4\alpha+1}}T^{4\alpha+1})$ $\begin{aligned} &e^{T(\frac{21}{2}+128\mathsf{K}^{4}_{\text{Total}}(1+T^{4\alpha}))}\\ &\qquad\times(\mathbb{E}[|\widehat{Y}_{0}^{\text{EM}}|^{4}]+1024\mathsf{K}^{4}_{\text{Total}}T(1+\mathbb{E}[|\hat{\theta}|^{4}])(1+T^{4\alpha})+8(d^{2}+4d+4)T)\end{aligned}$
@@
$C_{\mathsf{EMose},2}$ $O(de^{T^{2\alpha+1}}T^{4\alpha+1}\widetilde{\varepsilon}_{\text{AL}})$ $\begin{aligned} 2(C_{\mathsf{EM},2}(T)+\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(16C_{\mathsf{EM},2}(T)+32(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})))+2d\end{aligned}$
@@
$C_{4}(T,\epsilon)$ $O(\sqrt{d}e^{T^{2\alpha+1}}T^{3\alpha+1}\widetilde{\varepsilon}^{1/2}_{\text{AL}})$ $\begin{aligned} &\zeta^{-1/2}(T-\epsilon)^{1/2}e^{(1+(3/2)\zeta+2\mathsf{K}_{3}(1+2T^{\alpha}))(T-\epsilon)}\\ &\hskip 16.38895pt\times(C^{1/2}_{\mathsf{EMose},2}(1+2\mathsf{K}_{3}(1+2T^{\alpha}))+2\sqrt{2}\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2})\end{aligned}$
@@
$C_{\mathsf{EMose},4}$ $O(d^{2}e^{T^{4\alpha+1}}T^{8\alpha+1})$ $\begin{aligned} 8(C_{\mathsf{EM},4}(T)+\mathsf{K}^{4}_{\text{Total}}(1+T^{4\alpha})(1024C_{\mathsf{EM},4}(T)+8192(1+\mathbb{E}[|\hat{\theta}|^{4}])))+144d^{2}\end{aligned}$
@@
$\widetilde{C}_{4}(T,\epsilon)$ $O(de^{T^{4\alpha+1}}T^{4\alpha+1}\widetilde{\varepsilon}^{1/4}_{\text{AL}})$ $\begin{aligned} &\sqrt{2}e^{2(1+\zeta+\mathsf{K}_{3}(1+2T^{\alpha}+4\mathsf{K}_{3}(1+4T^{2\alpha})))(T-\epsilon)}\sqrt{T-\epsilon}\\ &\qquad\ \times\Bigg(\mathsf{K}_{4}^{2}\zeta^{-1}(1+4T^{2\alpha})C_{\mathsf{EMose},4}+4d(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\hskip 16.38895pt\qquad+2\zeta^{-1}\mathsf{K}_{1}^{2}(1+8(\widetilde{\varepsilon}_{\text{AL}}+|\theta^{*}|^{2}))\\ &\hskip 16.38895pt\qquad+4\zeta^{-1}d(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\hskip 16.38895pt\qquad\qquad\times[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))C_{\mathsf{EM},2}(T)\\ &\hskip 16.38895pt\hskip 16.38895pt\hskip 16.38895pt+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})]\\ &\hskip 16.38895pt\qquad+2[(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}C^{1/2}_{\mathsf{EMose},2}+2\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}]\\ &\hskip 16.38895pt\hskip 16.38895pt\times[d\sqrt{2}(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}]\Bigg)^{1/2}\end{aligned}$
$\epsilon_{\delta}$ - $\delta^{2}/(64(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{d})^{2})$
@@
$T_{\delta}$ - Obtained solving $T>T_{\delta}$ using Proposition 16, i.e., $\begin{aligned} &\ln(\mu(e^{2T}-1)+1)+(K/\mu+1)/(\mu(e^{2T}-1)+1)\\ &>\ln(4\sqrt{2}((\mathbb{E}[|X_{0}|^{2}])^{1/2}+\sqrt{d})/\delta)+2\int_{0}^{\epsilon}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t+K/\mu+1-\epsilon\end{aligned}$
@@
$\varepsilon_{\text{SN},\delta}$ - $\begin{aligned} (\delta^{2}\zeta/32)e^{-2(1+\zeta)(T-\epsilon)+4\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t}\end{aligned}$
@@
$\gamma_{\delta}$ - $\begin{aligned} &(\delta^{2}\zeta/16)(T-\epsilon)^{-1}e^{-2(1+(3/2)\zeta+2\mathsf{K}_{3}(1+2T^{\alpha}))(T-\epsilon)}\\ &\qquad\times(C^{1/2}_{\mathsf{EMose},2}(1+2\mathsf{K}_{3}(1+2T^{\alpha}))+2\sqrt{2}\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2})^{-2}\end{aligned}$
@@
$\widetilde{\gamma}_{\delta}$ - $\begin{aligned} \min\Bigg\{&(\delta/(4\sqrt{2}))^{1/\alpha}(T-\epsilon)^{-1/(2\alpha)}e^{-(2/\alpha)(1+\zeta+\mathsf{K}_{3}(1+2T^{\alpha}+4\mathsf{K}_{3}(1+4T^{2\alpha})))(T-\epsilon)}\\ &\times\Bigg(\mathsf{K}_{4}^{2}\zeta^{-1}(1+4T^{2\alpha})C_{\mathsf{EMose},4}+4d(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\hskip 16.38895pt+2\zeta^{-1}\mathsf{K}_{1}^{2}(1+8(\widetilde{\varepsilon}_{\text{AL}}+|\theta^{*}|^{2}))\\ &\hskip 16.38895pt+4\zeta^{-1}d(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\hskip 16.38895pt\qquad\times[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))C_{\mathsf{EM},2}(T)\\ &\hskip 16.38895pt\hskip 16.38895pt+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})]\\ &\hskip 16.38895pt+2[(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}C^{1/2}_{\mathsf{EMose},2}+2\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}]\\ &\hskip 16.38895pt\hskip 16.38895pt\times[d\sqrt{2}(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}]\Bigg)^{-1/(2\alpha)},1\Bigg\}\end{aligned}$
Table 4 provides the main notation used throughout this work and indicates where each symbol is introduced.
Table 4: List of the main notation.
Symbol Reference in the text
$\pi_{\mathsf{D}}$ Equation 1
$\hat{\theta}$ Assumption 1
$\widetilde{\varepsilon}_{\text{AL}}$ Assumption 1
$\partial U$ Definition 3
$K$ Assumption 2-(ii)
$\mu$ Assumption 2-(iii)
$\alpha$ Assumption 3.a
$\varepsilon_{\text{SN}}$ Assumption 4
$\kappa_{U}$ Definition 11
$f_{L}$ Equation 13
$\beta^{\text{OS},K,\mu}_{t}$ Equation 22
$B(t,0,\mu,K)$ Equation 23